Consider the following cumulative distribution function for X. 7 0.1 08 0.9 1.0 Fo) 0.3 0.6 (i) Determine the probability distribution. ii) Find P(X < 1). iii Find P(0 <XS5).
Answer #1

(i)

F(x) changes value at x = - 3, -1, 1, 3, 5, 7

So,

Range of X is given by:
{-3, -1, 1, 3, 5, 7}

So,

f(-3) = 0.1

f(-1) = F(-1) - F(-3) = 03 - 0.1 = 0.2

f(1) =F(1) - F(-1) = 0.6 - 0.3 = 0.3

f(3) = F(3) - F(1) = 0.8 - 0.6 = 0.2

f(5) = F(5) - F(3) = 0.9 - 0.8 = 0.1

f(7) = F(7) - F(5) = 1.0 - 0.9 = 0.1

Thus,

the probability distribution of x is given by:

X f(X)
- 3 0.1
- 1 0.2
1 0.3
3 0.2
5 0.1
7 0.1

(ii)

P(X<1) = f(-1) + f(-3) = 0.2 + 0.1 = 0.3

(iii)

P(0<X\leq5) = f(1) + f(3) + f(5) = 0.3 + 0.2 + 0.1 = 0.6